OpenStudy (ilovebmth1234):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)?

OpenStudy (ilovebmth1234):

x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

alones (alones):

Those have exponents?

OpenStudy (ilovebmth1234):

yes lol

OpenStudy (ilovebmth1234):

sorry the numbers next to x and y are exponents ^2

alones (alones):

Ugh I'm still confused

OpenStudy (ilovebmth1234):

x^2 + y^2 − 4x + 2y + 1 = 0 x^2 + y^2 + 4x − 2y + 1 = 0 x^2 + y^2 + 4x − 2y + 9 = 0 x^2 − y^2 + 2x + y + 1 = 0

OpenStudy (ilovebmth1234):

the sign ^ means the number after it is an exponent

OpenStudy (jskhupmang):

im pretty sure the answer should be the first two rows

OpenStudy (ilovebmth1234):

it can only be one

OpenStudy (jskhupmang):

I meant the first 3 row actually because a circle is x^2+y^2=r^2

OpenStudy (jskhupmang):

right

OpenStudy (ilovebmth1234):

it can only be one of these

OpenStudy (ilovebmth1234):

A) x2 + y2 − 4x + 2y + 1 = 0 B) x2 + y2 + 4x − 2y + 1 = 0 C) x2 + y2 + 4x − 2y + 9 = 0 D) x2 − y2 + 2x + y + 1 = 0

OpenStudy (jskhupmang):

like i said the center is (h,k) so in this case it is (-2,1)

OpenStudy (jskhupmang):

ok so basically you have to separate x and y in groups

OpenStudy (jskhupmang):

but in this case you have to work backward

Directrix (directrix):

The distance between these two points is the radius of the circle. center at (-2, 1) and passing through (-4, 1)? The points are on the same horizontal line and are 2 units apart.